### v1

# Rethinking Visual Layer Selection in Multimodal LLMs

Abstract
Multimodal large language models (MLLMs) have achieved impressive performance across a wide range of tasks, typically using CLIP-ViT as their visual encoder due to its strong text-image alignment capabilities. While prior studies suggest that different CLIP-ViT layers capture different types of information—shallower layers focusing on fine visual details and deeper layers aligning more closely with textual semantics—most MLLMs still select visual features based on empirical heuristics rather than systematic analysis. In this work, we propose a Layer-wise Representation Similarity approach to group CLIP-ViT layers with similar behaviors into shallow, middle, and deep categories and assess their impact on MLLM performance. Building on this foundation, we revisit the visual layer selection problem in MLLMs at scale, training LLaVA-style models ranging from 1.4B to 7B parameters. Through extensive experiments across 10 datasets and 4 tasks, we find that: (1) deep layers are essential for OCR tasks; (2) shallow and middle layers substantially outperform deep layers on reasoning tasks involving counting, positioning, and object localization; (3) a lightweight fusion of features across shallow, middle, and deep layers consistently outperforms specialized fusion baselines and single-layer selections, achieving gains on 9 out of 10 datasets. Our work offers the first principled study of visual layer selection in MLLMs, laying the groundwork for deeper investigations into visual representation learning for MLLMs.
1 Introduction
Recent advancements in Large Language Models (LLMs) [38, 39, 40, 48, 3] have catalyzed the development of Multimodal Large Language Models (MLLMs) [12, 3, 5], which aim to unify visual understanding with language reasoning – a capability foundational to applications ranging from robotic navigation to medical diagnosis [12, 2, 6, 3, 46]. While the integration of advanced LLMs has significantly enhanced the linguistic prowess of MLLMs, the vision component remains a critical yet underexplored bottleneck. Specifically, the selection of optimal visual features from hierarchical layers of Vision Transformers (ViTs) is often empirically driven [49, 16, 47] , lacking systematic justification.
Current MLLMs predominantly adopt a deep-layer bias: models like LLaVA [27, 28, 26] extract visual features from the penultimate layer of CLIP-ViT [40], while InternVL [6] variably uses the final or fourth-to-last layer, e.g. versions like InternVL 6B v1.2 and v1.5 use the final layer, while other variants within its series opt for the fourth-to-last layer as their visual representation. These choices, however, are rarely grounded in rigorous analysis. Prior works [10, 49, 16, 47, 4, 5] have demonstrated that ViT layers encode distinct semantic hierarchies (e.g., edges in shallow layers vs. object parts in deeper layers), yet their impact on MLLM performance remains unclear. This raises a pivotal question: Does the conventional reliance on deep layers truly maximize MLLM capabilities, or are we overlooking untapped potential in shallow/middle layers?
Although integrating features from multiple hierarchical levels has been extensively studied in computer vision [13, 25], its systematic exploration in the context of MLLMs is still lacking. While works like CLIP-InterpreT [10] reveal that different attention heads within ViT layers attend to distinct properties, there has been limited investigation into how such internal variations should inform the selection or fusion of visual features in multimodal models.
To address this gap, we propose Layer-wise Representation Similarity (LRS) to quantify the cosine similarity of hidden states across CLIP-ViT layers. Our analysis reveals three distinct visual representation spaces: shallow (layer 1-12), middle (layer 13-20), and deep (layer 21-24), as illustrated in Fig. 1.
Building upon this foundation, we first re-evaluate the effectiveness of different deep layers. Although the penultimate layer does not always yield the absolute best performance, we find that it consistently outperforms other deep layers across a range of LLM scales (1.4B, 2.7B, and 7B parameters). This superiority stems from its ability to preserve more visual details while maintaining strong text alignment, validating its widespread use. However, we also observe that as LLM size increases, gains diminish when visual quality becomes the bottleneck—particularly in hallucination-prone tasks. This highlights an important limitation: the final-layer degradation caused by CLIP’s contrastive training cannot be compensated for simply by scaling up the LLM. Thus, among deep layers, the penultimate layer remains the best default choice.
Having established the penultimate layer’s dominance within the deep space, we turn to an even more critical question: could earlier layers offer complementary strengths? Surprisingly, our experiments reveal that shallow and middle layers outperform deep layers on approximately one-third of the MME [9] sub-tasks, particularly those involving counting and positioning. Consistent patterns are observed in MMVet [51]. As shown in Tab. 1, representations from the layer 18 surpass those from the penultimate layer by 3% on CVBench [46]. Fig. 2 shows that, although shallow layers have the worst overall performance compared to the middle and deep layers, it still outperforms them in a notable number of sub-tasks. To validate the robustness of these findings, we evaluate performance across three data scales (665k, 737k, and 1M samples). Despite some task-specific variations, the core conclusion remains: shallow and middle layers harbor untapped potential.
Motivated by these insights, we propose a minimalist fusion strategy that integrates features from shallow, middle, and deep layers. Our design employs only a single linear layer, incurring minimal computational overhead while substantially improving visual representations. This simple yet effective approach demonstrates that lightweight fusion strategies can systematically unlock performance gains, offering a principled alternative to ad-hoc layer selection in future MLLM research. Although previous studies [49, 12, 7, 4] have explored hierarchical feature fusion or LLM-aligned feature selection, they often rely on empirical heuristics without a dedicated analysis of the intrinsic layer properties themselves. In contrast, our work is the first to systematically analyze information variation across ViT layers and how it aligns with LLMs under controlled experimental setups.
The key contributions of this paper can be summarized as follows:
To analyze visual layer relationships in CLIP-ViT, we propose LRS, identifying three distinct groups: shallow (1–12 layers), middle (13–20 layers), and deep (21–24 layers). (Sec. 3.3)
- (2)
- (3)
2 Related Work
2.1 Multimodal LLMs
The success of large language models (LLMs) [39, 48, 8, 3] and pretrained visual models [45, 40, 52] in their respective fields have laid a solid foundation for the emergence of multimodal LLMs. For instance, LLMs such as the LLaMA [8] series demonstrate powerful natural language processing capabilities, while CLIP-ViT [40] align the visual and textual modalities effectively through large-scale image-text contrastive learning.
Thanks to the contributions of numerous open-source language models, the parameters of LLMs [8] have expanded from 7B to 110B, significantly boosting multimodal performance. Additionally, some studies have focused on improving the visual capabilities of these models, with methods like dynamic splitting being employed to better understand high-resolution images.
Inspired by recent advances in computer vision, such as DenseNet [13] and FPN [25], some studies have explored integrating shallow and deep features within the ViT of multimodal language models. However, these approaches implement feature integration without a detailed analysis of layer-wise differences in multimodal tasks.
2.2 Vision Encoder in MLLMs
Serving as the “eyes” of Multimodal Large Language Models (MLLMs), the vision encoder sets the upper limit of the model’s perceptual capabilities. If certain information is absent from the representation, theoretically, the model cannot correctly answer questions that rely on this information as prior knowledge. CLIP, through image-text contrastive learning effectively aligns visual representation with text space and is widely adopted as the visual encoder in models such as LLaVA [27, 26], Qwen-VL [3], Flamingo [2], and BLIP [23]. Other foundational vision models, such as DINOv2 [37], SigLIP [52], ConvNeXT [32], are also utilized to build MLLMs. In this paper, we selected the widely used CLIP-ViT model as the focus of our layer-wise analysis.
2.3 Visual Fusion Methods
Visual feature fusion methods can be distinguished by whether visual tokens interact directly with text tokens. One approach [49, 47] merges hierarchical visual features through element-wise or channel-wise fusion before feeding them into the connector module, rather than directly using a single-layer ViT.
Another approach [2, 12, 5] uses a cross-attention mechanism to enable interaction between visual and textual representations, resulting in stronger text perception and comprehension capabilities.
However, understanding the differences in visual representations generated by different encoders remains a key research question in multi-modal large models. This study aims to provide a comprehensive analysis of layer-wise visual representations in MLLMs, offering valuable insights for the development of future visual fusion methods.
3 Model, Benchmark, and Analysis Method
3.1 Base Model
MLLMs have various structural designs, and our analysis of the role of visual layers is based on an LLaVA-like model. In this section, we will provide a detailed illustration of the structure of LLaVA-like models and the details of our implementation process.
LLaVA-like MLLM Structure
As one of the leading architectures in modern MLLMs, LLaVA consists of a CLIP-based visual encoder, a projection layer, and a pre-trained LLM. The encoder extracts image features and maps them to the LLM’s token space, enabling seamless vision-language interaction.
To explore the role of different visual layers in MLLMs, an ideal approach is to input each layer into a full MLLM and evaluate its performance on multimodal benchmarks. Given our focus on the visual side, a more efficient alternative is to use an LLaVA-like structure with a smaller LLM for layer-wise experiments.
In this equation, represents the initial hidden state of the model, which is derived from the input image through an embedding function . The input denotes a image with height , width , and channels.
Each hidden state is recursively computed using a TransformerBlock applied to , where and is the number of layers. Each TransformerBlock consists of multi-head self-attention (MSA) and feed-forward (FFN) layers, progressively refining feature representations. Our study focuses on analyzing these hidden states, termed visual representations, in the context of multimodal tasks.
Here, denotes the visual embedding, and denotes the text embedding. Each hidden state captures a different level of semantic information, with deeper layers generally encoding more abstract, high-level features. In our layer-wise experiment, is the input of the connector in MLLMs. Each visual representation space demonstrates unique characteristics across various tasks. In Sec 4, we provide a comprehensive analysis of our findings.
Implementation Details
We employ CLIP ViT-L/14 (336px) [40] as the visual encoder and 1.4B MobileLLaMA as the language model for efficiency analysis, with a one-layer MLP serving as the connector. Training follows a two-phase strategy aligned with LLaVA [27]. AdamW optimizer with a cosine annealing scheduler is used, with learning rates of 1e-3 (phase one) and 2e-5 (phase two), and batch sizes of 256 and 128. Training on four NVIDIA A100 80GB GPUs takes 2 hours for phase one and 8 hours for phase two. We adopt the LLaVA 1.5 [27] dataset, comprising 558K image-caption pairs for pre-training and 665K conversational instances for instruction tuning. Unless explicitly noted, the experimental setup remains the same.
| Layers | General | OCR | Vision-Centric | Hallu | ||||||||||
| MMEP | MMEC | MMB | SEEDB | GQA | TVQA | OCRB | CVB | CVB2D | CVB3D | RWD | MMVet | RefCOCO | POPE | |
| 1 | 750.1 | 211.4 | 0 | 25.30 | 40.55 | 7.99 | 24 | 40.14 | 34.87 | 45.42 | 38.17 | 9.9 | 5.73 | 70.21 |
| 2 | 790.2 | 212.5 | 0.34 | 25.76 | 41.24 | 7.96 | 23 | 40.21 | 34.34 | 46.08 | 37.12 | 10.1 | 5.36 | 71.01 |
| 3 | 742.7 | 219.2 | 0.17 | 25.00 | 41.76 | 8.10 | 28 | 42.69 | 37.89 | 47.50 | 36.08 | 10.4 | 6.87 | 72.33 |
| 4 | 788.4 | 239.6 | 0 | 25.53 | 42.26 | 8.50 | 23 | 42.69 | 36.71 | 48.67 | 37.25 | 10.2 | 7.61 | 72.34 |
| 5 | 813.2 | 220.7 | 0.17 | 25.26 | 42.74 | 8.22 | 21 | 41.30 | 33.69 | 48.92 | 36.86 | 10.9 | 8.25 | 72.91 |
| 6 | 838.8 | 227.8 | 0 | 25.23 | 43.16 | 8.26 | 24 | 41.69 | 35.97 | 47.42 | 36.86 | 11.5 | 9.31 | 75.18 |
| 7 | 815.6 | 235.7 | 0 | 25.74 | 44.90 | 8.75 | 25 | 43.02 | 37.12 | 48.92 | 37.39 | 10.6 | 11.10 | 75.44 |
| 8 | 857.7 | 237.5 | 0 | 25.48 | 46.14 | 8.77 | 25 | 41.43 | 36.85 | 46.00 | 36.99 | 11.2 | 10.79 | 76.20 |
| 9 | 889.7 | 232.8 | 0.17 | 27.72 | 47.02 | 9.05 | 28 | 40.53 | 36.23 | 44.83 | 37.12 | 13.0 | 10.06 | 77.84 |
| 10 | 903.4 | 228.2 | 0.17 | 26.61 | 48.39 | 9.03 | 30 | 41.8 | 36.19 | 47.42 | 37.39 | 11.4 | 13.50 | 77.46 |
| 11 | 935.3 | 224.3 | 0.52 | 26.58 | 49.85 | 10.65 | 32 | 42.51 | 37.27 | 47.75 | 36.86 | 14.2 | 12.14 | 79.24 |
| 12 | 980.1 | 232.1 | 0.09 | 26.85 | 50.39 | 16.58 | 70 | 41.81 | 36.7 | 46.92 | 38.56 | 12.6 | 11.20 | 80.63 |
| 13 | 964.0 | 252.5 | 0.09 | 26.33 | 51.14 | 18.12 | 91 | 41.71 | 35.75 | 47.67 | 37.25 | 11.6 | 12.50 | 81.39 |
| 14 | 984.2 | 265.4 | 0.69 | 34.07 | 51.83 | 22.86 | 130 | 42.69 | 36.12 | 49.25 | 39.08 | 13.8 | 14.37 | 81.97 |
| 15 | 1042.8 | 227.5 | 0.17 | 28.98 | 52.89 | 25.79 | 155 | 43.85 | 36.37 | 51.33 | 37.12 | 13.6 | 13.77 | 83.11 |
| 16 | 1069.5 | 225.4 | 0 | 27.95 | 52.81 | 28.08 | 166 | 43.26 | 36.61 | 49.92 | 38.04 | 13.7 | 15.41 | 84.32 |
| 17 | 1074.8 | 230.4 | 0.26 | 32.81 | 53.86 | 28.25 | 200 | 47.26 | 39.43 | 55.08 | 39.22 | 15.4 | 18.49 | 84.46 |
| 18 | 1088.7 | 237.1 | 29.38 | 52.06 | 54.37 | 31.44 | 200 | 47.29 | 41.17 | 53.42 | 39.48 | 14.3 | 17.04 | 84.26 |
| 19 | 945.1 | 236.8 | 20.02 | 44.64 | 48.32 | 18.27 | 121 | 45.69 | 37.21 | 54.17 | 35.95 | 13.2 | 18.22 | 81.47 |
| 20 | 1118.2 | 232.1 | 26.03 | 51.72 | 54.83 | 32.05 | 211 | 47.29 | 40.32 | 54.25 | 38.82 | 16.3 | 18.49 | 84.76 |
| 21 | 1041.4 | 212.5 | 0.95 | 35.42 | 49.47 | 28.10 | 190 | 44.37 | 39.32 | 49.42 | 39.87 | 14.5 | 17.09 | 81.91 |
| 22 | 1123.6 | 238.9 | 23.28 | 49.60 | 54.52 | 30.84 | 211 | 44.37 | 36.73 | 52.00 | 39.87 | 17.3 | 16.32 | 84.79 |
| 23 | 1142.7 | 245.0 | 35.31 | 52.84 | 54.61 | 33.73 | 233 | 44.26 | 38.02 | 50.50 | 45.36 | 18.0 | 17.08 | 84.00 |
| 24 | 1114.1 | 243.5 | 32.65 | 51.09 | 53.61 | 30.63 | 197 | 46.68 | 39.78 | 53.58 | 43.92 | 16.1 | 17.08 | 83.65 |
3.2 Evaluation Benchmarks
To comprehensively explore and evaluate various visual representations, we classified the benchmarks into four categories following previous work [46]: (1) General tasks, (2) OCR (Optical Character Recognition) tasks, (3) Vision-centric tasks, and (4) Hallucination tasks.
The General tasks category includes MME [9] (yes/no questions on attributes like existence and color), MMBench [30] (multiple-choice across diverse aspects), SEEDBench [22] (spatial and temporal reasoning), and GQA [15] (complex real-world VQA). The OCR category features TextVQA [44] and OCRBench [31], both assessing OCR capabilities. The Vision-centric category includes CVBench [46] (evaluating spatial relations and depth), RealWorldQA (real-world QA), MMVet [51] (general multimodal assessment), and RefCOCO [50] (visual grounding). Finally, the Hallucination category includes POPE [24], which evaluates object hallucination.
3.3 Layer-wise Representation Similarity
To examine the behavioral patterns of different visual layers, we propose an analysis method named Layer-wise Representation Similarity (LRS) to analyze the relationships between various visual layers.
Specifically, we compute the cosine similarity between hidden states of ViT layers, where values closer to indicate higher similarity. Fig. 1(a) shows the average cosine similarity of different visual layers. The formalized division process is detailed in the following equation:
We define the hidden state matrix of CLIP-ViT as:
where is the number of layers, and each row represents the hidden state of the -th layer. The cosine similarity matrix is computed as:
where computes the pairwise dot products between layer hidden states. The denominator normalizes the values using the -norm, defined as:
ensuring that values lie within . The absolute value guarantees that all elements are within the range . Then, we compute the average of across four tasks.
From Fig.1(a), we can identify three distinct representation spaces. As illustrated in Fig.1(b), shallow layers contribute minimally to OCR tasks, performance improves rapidly in the middle layers, and reaches its peak in the deep layers. This pattern suggests that visual layers within the same partition exhibit similar behaviors.
Based on behavioral similarity, we categorize the 24 CLIP-ViT visual layers into three groups: shallow layers (1 to 12), middle layers (13 to 20), and deep layers (21 to 24).
4 Experiment: Layer-wise Exploration
In this section, we address Q1: Can shallow and middle visual representations outperform deep one? We conduct a layer-wise analysis and evaluate across 10 benchmarks to compare different representations.
4.1 Settings
Layer-wise Analysis
To fully explore each layer’s potential, we input representations from each layer of the CLIP-ViT model into the connector. We conduct a systematic evaluation of visual representations across multimodal benchmarks. Specifically, we take through from Sec. 3 as inputs, followed by two-stage training. This setup helps assess the upper bound of visual representations in multimodal tasks.
4.2 Exp-I: Effect of Shallow and Middle Layers
Potential of Shallow and Middle layers
We first conduct an investigation into the shallow and middle representation spaces. The performance of these two spaces as shown in Tab. 1, several key insights emerge from the results:
The middle layer has the potential to perform best: Although the middle layer’s information has not been fully processed, it still achieves the best performance on one-third of the benchmarks. Specifically, compared to the penultimate layer, layer 14 achieves a 20-point higher score on MME-Cognitive, layer 18 outperforms by 3% on CVBench, layer 17 surpasses by 1.4% on RefCOCO, and layer 20 exceeds by 0.2% on GQA. This demonstrates the potential of middle layers.
The middle layers generally perform better on vision-centric tasks: Fig. 3(a) illustrates the performance of different layers across subtasks in the MME dataset, showing that position and existence tasks benefit more from middle layer representations. As depicted in Fig. 3(b), the penultimate layer achieves top performance in only three out of eleven subtasks, whereas the shallow and middle layers yield optimal results in seven. Similarly, in Fig. 3(c), one-third of the best performing results, such as those in spatial relations, physical relations, and cross fine-grained perception originate from the shallow and middle layers. A comparable trend is observed in SEEDBench (Fig. 3(d)), where middle layers produce optimal results in nearly half of the subtasks, including Instance Attribute, Instance Location, Instance Interaction, and Text Recognition.
Hallucination issues can be effectively handled at the middle layers: As shown in Tab. 1, POPE results indicate that hallucination issues are most prominent in the shallow representation space, with minimal variation between the middle and deep layers. Notably, in the middle representation space, half of the layers outperform the penultimate layer on the POPE. This is because, comparatively, the challenge of the task lies more in perception than comprehension. In Sec. 5.3, we provide a detailed analysis showing that further experiments with larger LLMs consistently support this finding.
4.3 Exp-II: Effect of Deep Layer
Which layer performs the best in deep space?
A widely accepted but rarely questioned assumption is which layer of visual representation should be used as input for MLLMs. Our findings reveal that utilizing the final layer visual representation as the input is not the optimal choice and may not even be a close second:
The last layer is not the optimal choice: As shown in Fig. 2, the last layer does not perform the best on any benchmark. Additional experimental results in Sec. 6.2 further validate this observation. The underlying reason lies in the CLIP model’s training mechanism, where supervision primarily focuses on aligning the final layer token with text embeddings. The token in the final layer is optimized by CLIP’s contrastive loss, making it highly specialized for the image-text matching task. However, this optimization process, driven by the attention mechanism, most significantly suppresses local details in the last layer.
Penultimate layer as the optimal choice: As shown in Fig. 2, the penultimate layer consistently achieves the best results. This is because it preserves more visual information while maintaining a text alignment capability second only to the final layer. This balance represents the optimal trade-off between visual richness and textual alignment.
Deep Layers are Essential for OCR Tasks
As shown in Fig. 1, shallow layers provide negligible text information for the LLM. A clear boundary exists between the shallow and middle layers, with layer 12 marking the transition point. Layers before this point fail to contribute meaningfully to text processing, while a sharp performance gain occurs immediately afterward. This can be attributed to two essential requirements for OCR tasks:
Rich fine-grained visual features in visual representation: In OCR tasks, a strong perception of details is often required. Therefore, these fine-grained details must be embedded in the representation with sufficiently strong signals to be effectively utilized by the LLM.
The visual representation has already aligned with the text space: Despite containing rich visual details, shallow layers lack intrinsic alignment with textual representations, limiting their usefulness for OCR tasks. Fig. 1(a) confirms this with notably low cosine similarity between shallow and deep layers, highlighting their distinct feature representations. This discrepancy poses a challenge for the connector, which can only align features originating from the deep (text) space or adjacent middle layers.
| Data Scale | Layers | General | OCR | Vision-Centric | Hallu | ||||||||
| MMEP | MMEC | MMB | SEEDB | GQA | TVQA | OCRB | CVB | CVB2D | CVB3D | RWQA | POPE | ||
| 665k | 3 | 742.7 | 219.3 | 0.17 | 25.00 | 41.76 | 8.10 | 28 | 42.69 | 37.89 | 47.50 | 36.08 | 72.33 |
| 18 | 1088.8 | 237.1 | 29.38 | 52.06 | 54.37 | 31.44 | 200 | 47.29 | 41.17 | 53.42 | 39.48 | 84.26 | |
| 23 | 1142.8 | 245.0 | 35.31 | 52.84 | 54.61 | 33.73 | 233 | 44.26 | 38.02 | 50.50 | 45.36 | 84.00 | |
| 24 | 1114.1 | 243.6 | 32.65 | 51.09 | 53.61 | 30.63 | 197 | 46.68 | 39.78 | 53.58 | 43.92 | 83.65 | |
| 737k | 3 | 845.9 | 225.4 | 0.26 | 26.33 | 44.12 | 8.34 | 27 | 42.36 | 35.81 | 48.92 | 37.52 | 74.98 |
| 18 | 1093.1 | 226.4 | 43.04 | 56.33 | 56.98 | 35.98 | 270 | 48.87 | 46.57 | 51.17 | 43.40 | 86.18 | |
| 23 | 1163.7 | 230.0 | 48.37 | 55.55 | 56.77 | 36.41 | 265 | 48.09 | 47.59 | 48.58 | 41.83 | 86.22 | |
| 24 | 1121.7 | 258.6 | 46.05 | 55.43 | 56.34 | 36.09 | 255 | 43.63 | 38.01 | 49.25 | 44.44 | 85.09 | |
| 1M | 3 | 871.8 | 215.4 | 13.40 | 40.21 | 45.67 | 8.03 | 26 | 43.04 | 37.74 | 48.33 | 39.08 | 74.08 |
| 18 | 1145.9 | 213.2 | 42.44 | 56.72 | 57.74 | 35.68 | 267 | 54.08 | 56.33 | 51.83 | 43.14 | 84.06 | |
| 23 | 1214.4 | 249.3 | 52.92 | 58.58 | 57.91 | 37.24 | 263 | 53.48 | 53.96 | 53.00 | 43.27 | 84.03 | |
| 24 | 1192.0 | 245.0 | 47.34 | 57.62 | 57.21 | 36.45 | 264 | 47.88 | 42.84 | 52.92 | 44.97 | 84.58 | |
5 Experiment: Effect of Data and Model Scale
To answer Q2: If the answer to Q1 is ‘yes’, does it still hold under different data and model scales? We extended our model on LLaVA-like structure but variants with different model sizes and using different training dataset scales.
5.1 Settings
More Training Data
Motivated by recent work [53], we aim to investigate the impact of training data scales. We adopt the same training procedure detailed in Sec. 4.1. In the first stage, we use the LLaVA 558k dataset [27]. For the second stage, we evaluate three dataset configurations: (1) LLaVA 665K, (2) Cambrian-1 737K [46], an expansion of the 665K dataset with additional OCR data, and (3) a custom 1M dataset that builds on the 737K dataset by incorporating data specifically curated for Counting and Spatial Relation tasks. The data for these tasks is sourced from CLEVR [17] and TallyQA [1]. The detailed composition of the dataset can be found in Appendix 9.
Different LLM Sizes
Building on the original 1.4B experiments, we extend our study to include MobileLLaMA [7] 2.7B and Vicuna v1.5 7B. These LLMs are selected due to their similar architectures, making them well-suited for investigating the impact of different LLM sizes.
LRS-based layer selection
Due to computational constraints, we do not conduct a full layer-wise analysis across different data scales. Instead, leveraging insights from Sec 4.2, we select representative layers from the shallow (layer 3), middle (layer 18), and deep (layers 23 and 24) representation spaces to examine how variations in data scale affect model performance.
5.2 Exp-I: Effect of Data Scales
As shown in Tab. 2, our findings indicate that the key conclusions in Sec. 4.2 remain valid across different training data scales. As data scales up, we uncover the following key insights:
Increasing data scale boosts the penultimate layer’s perception ability: After incorporating additional OCR and vision-centric data, the middle layers continue to outperform the penultimate layer, albeit with a decreasing gap. Since the visual encoder remains frozen during training, this suggests that compared to the middle layers, fine-grained information is less explicitly preserved in the deep layers. In other words, fine-grained details in the middle layers are more readily utilized by the LLM through the connector, whereas those in the deep layers are harder to extract, requiring larger amounts of data to activate the connector and enable the LLM to effectively capture these fine-grained features.
The penultimate layer does not always perform best, regardless of data scale: On the 665K dataset, layer 18 outperforms the penultimate layer by 3% on CVBench. After incorporating OCR training data and subsequently adding vision-centric data, it maintains its lead, though the margin narrows to 0.8% and 0.5%, respectively. on the 737k dataset, layer 18 occasionally outperforms penultimate layer, surpassing the penultimate layer by 0.8% in SEEDBench and 0.2% in GQA. Moreover, as shown in Appendix 8 and 9, results from just four layers across various subtasks already demonstrate that the penultimate layer does not consistently achieve the best performance, let alone when considering all 24 layers.
Increasing the data scale does not enhance the capability of shallow layers in OCR tasks: Layer 3 shows similar performance on OCR tasks before and after incorporating OCR training data, indicating that increasing task-specific data alone cannot overcome the inherent limitations of shallow representations. This is due to their natural misalignment with the textural feature space. In contrast, middle layers, though only partially aligned, still benefit from additional training.
Middle and deep layers exhibit minimal differences on hallucination task: Across all dataset sizes, shallow layers perform significantly worse than middle and deep layers. But, the performance difference between middle and deep layers is still negligible.
5.3 Exp-II: Effect of LLM Scales
In this section, we further validate that these findings hold across LLMs of different sizes. While conclusions from small models may not fully generalize to larger ones, the main findings still hold, as detailed below:
The penultimate Layer remains the optimal choice in deep space regardless of LLM size: As shown in Fig. 4, the penultimate layer consistently outperforms the final layer across LLMs of 1.4B, 2.7B, and 7B, reinforcing our findings. This indicates that CLIP-ViT’s final-layer visual degradation, driven by its training paradigm, cannot be offset by a stronger LLM. Furthermore, increasing LLM size does not yield significant improvements in POPE performance, indicating that hallucination bottlenecks in MLLMs stem primarily from the quality of the visual representation.
Aligning with the prior results regardless of LLM size: As shown in Appendix 6 and 7, under the 2.7B model, the penultimate layer fails to achieve the best performance in half of the MME subtasks. For instance, Layer 18 performs better in Count, Position, and Existence tasks. Similarly, in SEEDBench, the penultimate layer underperforms in more than half of the subtasks, such as Spatial Relation and Instance Interaction. A similar trend is observed in the 7B model.
6 Experiment: Visual Feature Fusion
| Models | General | OCR | Vision-Centric | Hallu | |||||||||
| MMEP | MMEC | MMB | SEEDB | GQA | TextVQA | OCRB | CVB | CVB2D | CVB3D | RWQA | POPE | Win | |
| Baseline() | 1142.8 | 245.0 | 35.31 | 52.84 | 52.84 | 33.73 | 233 | 44.26 | 38.02 | 50.50 | 45.36 | 84.00 | 9/10 |
| + 18 | 1148.5 | 228.9 | 46.91 | 57.01 | 56.80 | 37.66 | 273 | 44.73 | 35.79 | 53.67 | 45.49 | 84.51 | 8/10 |
| + 1+18 | 1155.4 | 246.8 | 48.54 | 56.75 | 56.68 | 36.53 | 236 | 45.65 | 36.21 | 55.08 | 46.93 | 84.56 | 7/10 |
| + 17+18 | 1182.5 | 220.7 | 48.80 | 56.68 | 56.48 | 38.29 | 263 | 45.38 | 36.25 | 54.50 | 44.71 | 85.50 | 6/10 |
| DC-STI | 1142.4 | 218.9 | 48.02 | 57.23 | 56.86 | 36.42 | 226 | 43.83 | 35.10 | 52.58 | 44.44 | 86.38 | 8/10 |
| 1166.5 | 241.8 | 48.71 | 57.26 | 56.61 | 36.70 | 241 | 43.19 | 34.96 | 51.42 | 44.44 | 84.45 | 7/10 | |
| DC-DCI | 1145.0 | 253.2 | 47.85 | 57.16 | 56.92 | 37.54 | 257 | 45.60 | 35.83 | 54.92 | 45.10 | 84.95 | 7/10 |
| MMFuser | 1149.5 | 238.9 | 49.65 | 56.21 | 56.59 | 35.43 | 245 | 45.70 | 36.89 | 54.50 | 44.83 | 84.53 | 8/10 |
| 1157.2 | 236.1 | 49.22 | 57.23 | 57.35 | 37.70 | 265 | 44.56 | 36.53 | 52.58 | 45.75 | 84.82 | - | |
In this section, we aim to address Q3: If Q1 and Q2 are affirmed, can we leverage the advantages of shallow and middle layers to design better MLLMs?
6.1 Settings
The equation below shows that the simplest visual feature fusion mechanism,
where denotes the set of selected layers, and each represents the feature representation extracted from layer with a dimension of . Here, is the number of visual tokens, and is the feature dimensionality of each token. The concatenation function merges these representations along the feature dimension, producing an output of size , where denotes the number of concatenated layers. The resulting is then fed into the Connector. Subsequently, we explore various fusion strategies to optimize feature integration.
6.2 Exp-I: Fusion Selection via LRS Analysis
The Layer-wise Representation Similarity (LRS) analysis reveals the potential of shallow and middle layers. Based on it, we selected representative layers from shallow, middle, and deep layers for fusion experiments, aiming to preliminarily explore the role of different representation spaces in fusion methods.
As shown in Tab. 3, we optimize our approach by selecting layers from different representation spaces for feature fusion. In the shallow layer, we select layers 1 and 3 as representatives. Layer 1, being the most chaotic, primarily captures early-stage visual features, while layer 3 is still in a chaotic state but performs relatively well. In the middle layer, we choose layers 18 and 17, as they achieve the first and second-best performance within this representation space. For the deep layer, we select layer 23, as it demonstrates the highest overall performance. The baseline configuration considers only layer 23 as the visual representation. We systematically explore different combinations of to analyze their impact:
1.
Incorporating a middle layer: .
2.
Incorporating middle, shallow layer: .
3.
Incorporating two middle layer: .
4.
Quantile-driven layer selection: .
5.
RLS-driven optimal combination: .
and our main findings can be summarized as follows:
Visual representation fusion improves performance: As demonstrated in Tab. 3, all fusion methods exhibit a substantial performance improvement over the baseline. The best fusion approach outperforms the baseline in 9 out of 10 benchmarks. These results suggest that a direct visual feature fusion strategy can effectively enhance performance across various task categories, including General, OCR, vision-centric, and Hallucination tasks.
On Layer Combinations: Compared to two-layer fusion (), three-layer fusion () results in more stable performance improvements. However, the inclusion of shallow features diminishes the effectiveness of fusion strategies in OCR tasks. Feature layer selection guided by LRS analysis consistently outperforms selection methods based on quantiles (). As demonstrated in Sec. 4.3, the representation space of shallow layers has a minimal impact on OCR tasks. Therefore, incorporating shallow layers into feature fusion may negatively affect performance.
6.3 Exp-II: Fusion Method Comparison
Under the same experimental conditions, we select different combinations of visual representations based on LRS analysis. Despite its simplicity, this approach consistently surpasses existing fusion strategies across multiple benchmarks, underscoring the significance of our findings.
We evaluate four state-of-the-art fusion methods, including three from DenseConnector (DC) [49] and one from MMFuser [4]. As shown in Tab. 3, our method outperforms STI, SCI, and DCI on 8, 7, and 7 benchmarks, respectively. Compared to MMFuser [4], our approach demonstrates superior performance on 8 benchmarks. These results highlight the significant potential of visual feature fusion strategies in enhancing MLLMs and offering guidance for developing future fusion strategies.
7 Conclusion
This study presents a comprehensive layer-wise analysis, revealing that shallow and middle representation spaces can surpass the performance of deep layers. Evaluations across diverse data and model scales further substantiate this finding. Furthermore, we introduce a straightforward yet highly effective fusion strategy for visual feature integration, delivering substantial improvements over the baseline. Our findings offer a foundation for advancing future research in fusion methodologies.
Supplementary Material
We provide some additional information as a supplementary material. This material is divided into three sections:
Detailed analysis of visual representations
Experiment Details
Guidelines on Alignment
8 Visual Representations
In this section, we explore the variations in visual representation spaces across four tasks. As illustrated in Figure 5, the partitioning of visual representations is minimally influenced by the nature of the tasks. In other words, the shallow, middle, and deep representation spaces exhibit remarkable stability, maintaining consistent structures across various tasks.
Figure 6 further highlights the distinct characteristics of different representation spaces through the results on TextVQA and OCRBench. These results clearly demonstrate that shallow layers are ineffective for OCR tasks, with performance progressively improving in the middle space and being adequately addressed only in the deep space.
Another example, as shown in Table 4, when the penultimate layer visual representations are replaced with those from other layers of the visual encoder, layers 20 to 24, belonging to the deep visual representation space, show no signs of catastrophic performance degradation.
| Layer | MME-P | MME-C | OCRB | TextVQA | RefCOCO |
| 24 | 1153.5 | 306.0 | 266 | 41.16 | 47.56 |
| 23 | 1509.9 | 365.3 | 314 | 46.10 | 49.04 |
| 22 | 1451.1 | 366.7 | 304 | 44.76 | 47.46 |
| 21 | 1368.8 | 293.2 | 287 | 41.59 | 40.47 |
| 20 | 1259.2 | 265.7 | 271 | 39.11 | 44.31 |
| 19 | 1183.3 | 267.1 | 240 | 36.76 | 42.83 |
| 18 | 1083.7 | 237.8 | 205 | 32.18 | 36.04 |
| 17 | 993.6 | 255.7 | 156 | 27.92 | 31.07 |
| 16 | 901.0 | 256.7 | 116 | 23.37 | 19.96 |
| 15 | 790.0 | 253.9 | 94 | 17.96 | 14.85 |
| Data | Size |
| LLaVA [29] | 158K |
| + ShareGPT [42] | 40K |
| + VQAv2 [11] | 83K |
| + GQA [14] | 72K |
| + OKVQA [34] | 9K |
| + OCRVQA [36] | 80K |
| + A-OKVQA [41] | 66K |
| + TextCaps [43] | 22K |
| + RefCOCO [19, 33] | 48K |
| + VG [21] | 86K |
| LLaVA-1.5 [26] | 665K |
| + AI2D [20] | 16K |
| + DocVQA [35] | 15K |
| + DVQA [18] | 13K |
| Cambrian-737k [46] | 737K |
| + CLEVR [17] | 215k |
| + TallyQA [1] | 77K |
| Customized-1M | 1M |
9 Experiment Details
9.1 Visualization of Fusion Performance
As shown in Fig. 7, our fusion strategy achieves highly competitive performance.
9.2 Composition of Three Scale Datasets
The following datasets are incorporated to enhance the model’s capabilities across multiple multimodal tasks:
AI2D (Allen Institute for AI Diagram Dataset) [20] AI2D is designed for visual reasoning and diagram understanding, featuring annotated diagrams with textual descriptions and Q&A pairs. It is particularly useful for multimodal reasoning and visual question answering (VQA) tasks.
DocVQA (Document Visual Question Answering) [35] DocVQA focuses on visual question answering over document images, where questions pertain to scanned documents, OCR-recognized text, and textual reasoning. This dataset is valuable for document comprehension, text recognition, and multimodal reasoning.
DVQA (Diagrammatic Visual Question Answering) [18] DVQA is designed for visual question answering over diagrams and charts, covering questions related to bar charts, pie charts, and scientific illustrations. It evaluates the model’s ability to read structured visual information and perform reasoning based on graphical representations.
CLEVR (Compositional Language and Elementary Visual Reasoning) [17] CLEVR is a synthetic dataset for visual reasoning, containing 3D-rendered scenes with structured questions that require reasoning based on attributes, object relationships, and compositional logic. It is widely used to assess a model’s capability in compositional and multi-step reasoning.
TallyQA [1] TallyQA is a dataset specifically designed for object counting tasks, where questions require the model to accurately count objects in an image. It evaluates the model’s ability to attend to relevant objects, integrate global and local information, and perform numerical reasoning.
9.3 Evaluation Metrics
We provide a comprehensive explanation of the evaluation methods, categorizing them into three distinct types based on the evaluation metrics:
For benchmarks such as MME-Perception, MME-Cognition, OCRBench, and MMVet, we adopt the common approach of directly using the dataset-defined scores. We follow this established approach to maintain consistency and comparability in evaluations.
Using Accuracy directly as the evaluation metric. This applies to benchmarks such as MMBench, SEEDBench, GQA, TextVQA, CVBench, RealworldQA, and POPE.
In evaluating the RefCOCO dataset, we use CIDEr (Consensus-based Image Description Evaluation) as the primary evaluation metric.
To facilitate evaluation, we use lmms-eval as our primary evaluation tool. For the MMVet dataset, evaluations must be conducted on the official platform by uploading the necessary data. Regarding the CVBench 3D tasks, where models generally exhibit weaker instruction-following performance, we employ the DeepSeek API as the judge. This tool provides results consistent with GPT-4o but is significantly more cost-effective.
9.4 The impact of LLMs
Additional experiments are conducted on different sizes of large language models to investigate their impact on visual information processing. We validate our conclusions on LLMs of 2.7b and 7b sizes. Due to computational resource constraints, we selected representative layers from the three representation spaces to conduct experiments on subtasks of MME and SEEDBench. As shown in Table 6, the penultimate layer does not consistently achieve the best performance on MME. The commonly used penultimate layer achieves optimal performance on 6 out of 14 subtasks, while other layers, such as Layers 3, 18, and 24, demonstrate superior performance on the remaining subtasks. This observation aligns with prior findings, suggesting that middle layers can exhibit superior performance over deeper layers on certain CV-centric tasks. Notably, layer 18 outperforms the penultimate layer in tasks such as Count, Position, and Existence.
| Model Size | 2.7b | 7b | ||||||
| Layers | 3 | 18 | 23 | 24 | 3 | 18 | 23 | 24 |
| Code Reasoning | 52.50 | 47.50 | 47.50 | 40.00 | 50.00 | 40.00 | 42.50 | 45.00 |
| Artwork | 53.00 | 65.00 | 65.75 | 64.50 | 50.00 | 69.25 | 71.00 | 70.75 |
| Celebrity | 46.76 | 49.12 | 64.12 | 58.82 | 51.76 | 59.41 | 74.71 | 74.41 |
| Numerical Calculation | 50.00 | 50.00 | 42.50 | 25.00 | 47.50 | 45.00 | 37.50 | 37.50 |
| Text Translation | 50.00 | 50.00 | 50.00 | 50.00 | 65.00 | 65.00 | 47.50 | 67.50 |
| Count | 50.00 | 65.00 | 61.67 | 58.33 | 56.67 | 85.00 | 85.00 | 80.00 |
| Color | 53.33 | 83.33 | 86.67 | 88.33 | 78.33 | 91.67 | 91.67 | 91.67 |
| Commonsense Reasoning | 52.86 | 60.71 | 64.29 | 62.14 | 57.86 | 69.29 | 73.57 | 72.86 |
| Position | 48.33 | 71.67 | 71.67 | 71.67 | 61.67 | 71.67 | 75.00 | 80.00 |
| OCR | 50.00 | 67.50 | 72.50 | 65.00 | 55.00 | 77.50 | 75.00 | 70.00 |
| Landmark | 59.50 | 75.25 | 80.25 | 76.75 | 66.25 | 78.00 | 86.00 | 84.50 |
| Scene | 73.75 | 87.75 | 87.75 | 89.00 | 80.00 | 85.50 | 85.50 | 87.50 |
| Existence | 83.33 | 98.33 | 96.67 | 95.00 | 81.67 | 96.67 | 96.67 | 95.00 |
| Posters | 37.41 | 59.18 | 65.65 | 64.29 | 51.02 | 74.83 | 81.63 | 83.33 |
| Model Size | 2.7b | 7b | ||||||
| Layers | 3 | 18 | 23 | 24 | 3 | 18 | 23 | 24 |
| Scene Understanding | 36.04 | 68.84 | 68.87 | 69.79 | 50.70 | 73.50 | 73.91 | 73.59 |
| Instance Identity | 32.33 | 62.26 | 62.92 | 62.59 | 41.40 | 67.78 | 70.29 | 70.56 |
| Instance Attribute | 40.35 | 62.19 | 59.07 | 60.46 | 50.48 | 69.09 | 68.70 | 68.21 |
| Instance Location | 37.53 | 52.25 | 49.69 | 53.78 | 43.46 | 61.04 | 61.45 | 59.71 |
| Instance Counting | 25.70 | 43.07 | 47.57 | 45.20 | 33.02 | 56.89 | 57.13 | 57.29 |
| Spatial Relation | 33.03 | 42.47 | 40.64 | 43.99 | 41.55 | 52.05 | 49.62 | 51.45 |
| Instance Interaction | 34.02 | 64.95 | 54.64 | 64.95 | 48.45 | 63.92 | 67.01 | 71.13 |
| Visual Reasoning | 35.05 | 67.07 | 72.51 | 72.21 | 53.47 | 75.23 | 78.85 | 77.04 |
| Text Recognition | 44.71 | 21.18 | 21.18 | 24.71 | 38.82 | 34.12 | 47.06 | 43.53 |
As illustrated in Table 7, while performance varies slightly across the subtasks of SEEDBench, the penultimate layer achieves the best performance on only 3 out of 9 subtasks. These results provide strong empirical evidence that shallow and middle layers can outperform deeper layers on specific subtasks.
9.5 The impact of Data Scale
The conclusion remains valid across different data scales. Under the 737k data scale, half of the subtasks in the MME dataset achieve optimal performance using the penultimate layer. However, for tasks like Count, Position, and Existence, the middle visual representation layer (Layer 18) demonstrates either superior or comparable performance. Similarly, results under the 1M data scale also show that half of the optimal performances are achieved on layers other than the penultimate one. The results for SEEDBench subtasks, as presented in Table 9, further support this observation. At the 737k data scale, Layer 18 from the middle representation space achieves the best performance on 5 out of 9 subtasks, while the penultimate layer excels in only 3 subtasks. Likewise, under the 1M data scale, half of the subtasks continue to achieve their best performance on layers other than the penultimate one. These findings consistently demonstrate across varying data scales that shallow and middle layers have the potential to outperform deep layers in certain scenarios.
| Data Scale | 737k | 1M | ||||||
| Layers | 3 | 18 | 23 | 24 | 3 | 18 | 23 | 24 |
| Code Reasoning | 50.00 | 47.50 | 45.00 | 42.50 | 47.50 | 47.50 | 45.00 | 47.50 |
| Artwork | 51.00 | 59.25 | 64.50 | 61.75 | 53.75 | 65.00 | 68.00 | 66.50 |
| Celebrity | 48.82 | 55.88 | 64.12 | 62.65 | 52.35 | 62.06 | 68.53 | 65.59 |
| Numerical Calculation | 47.50 | 37.50 | 35.00 | 47.50 | 50.00 | 30.00 | 45.00 | 47.50 |
| Text Translation | 50.00 | 50.00 | 50.00 | 50.00 | 50.00 | 50.00 | 50.00 | 50.00 |
| Count | 55.00 | 55.00 | 55.00 | 50.00 | 58.33 | 58.33 | 60.00 | 60.00 |
| Color | 60.00 | 75.00 | 76.67 | 75.00 | 61.67 | 78.33 | 80.00 | 80.00 |
| Commonsense Reasoning | 54.29 | 56.43 | 60.71 | 61.43 | 52.14 | 57.14 | 62.14 | 58.57 |
| Position | 51.67 | 70.00 | 63.33 | 70.00 | 50.00 | 70.00 | 73.33 | 70.00 |
| OCR | 50.00 | 55.00 | 57.50 | 52.50 | 55.00 | 52.50 | 55.00 | 55.00 |
| Landmark | 63.25 | 71.00 | 77.50 | 74.25 | 61.75 | 72.50 | 77.50 | 76.00 |
| Scene | 69.25 | 83.50 | 85.00 | 84.25 | 74.00 | 84.50 | 84.00 | 85.75 |
| Existence | 85.00 | 96.67 | 96.67 | 96.67 | 80.00 | 96.67 | 96.67 | 95.00 |
| Posters | 38.44 | 50.34 | 55.44 | 56.80 | 37.41 | 52.72 | 55.44 | 54.76 |
| Data Scale | 737k | 1M | ||||||
| Layers | 3 | 18 | 23 | 24 | 3 | 18 | 23 | 24 |
| Scene Understanding | 26.85 | 67.35 | 68.24 | 67.57 | 44.02 | 65.71 | 68.97 | 67.57 |
| Instance Identity | 23.65 | 56.36 | 55.60 | 55.87 | 35.06 | 57.89 | 61.82 | 58.77 |
| Instance Attribute | 26.03 | 59.93 | 58.36 | 57.69 | 44.27 | 61.93 | 61.45 | 60.59 |
| Instance Location | 26.69 | 50.20 | 47.75 | 48.26 | 39.37 | 48.67 | 49.69 | 49.08 |
| Instance Counting | 27.34 | 41.93 | 40.13 | 41.19 | 32.04 | 42.38 | 44.91 | 45.97 |
| Spatial Relation | 29.68 | 41.86 | 38.81 | 38.96 | 37.90 | 42.92 | 39.57 | 39.57 |
| Instance Interaction | 21.65 | 46.39 | 51.55 | 49.48 | 38.14 | 50.52 | 51.55 | 52.58 |
| Visual Reasoning | 27.79 | 62.54 | 67.98 | 69.18 | 43.20 | 60.42 | 68.88 | 67.07 |
| Text Recognition | 21.18 | 34.12 | 48.24 | 43.53 | 41.18 | 17.65 | 55.29 | 42.35 |
10 Guidelines on Alignment
What defines a good visual representation in multimodal models? Firstly, an ideal visual representation must simultaneously provide rich visual information and effectively align with textual modal. An MLLM can only correctly answer queries when the information corresponding to the given instruction is explicitly embedded within the visual representation. However, when the visual representation fails to deliver the necessary information, the model becomes prone to hallucination problems. Secondly, alignment with the textual modality is essential for a large language model to understand and process information from a different modality. This alignment ensures that the rich visual content is effectively leveraged. In a word, the CLIP series models currently offer the best trade-off between these two dimensions.